OpenStudy (anonymous):

Solve for x: x + 5/6 = 1/3 (okay so basically it's x plus 5 over 6 equals 1 over 3. A. 3 B. -3 C. 25/3 (25 over 3) D. 1

OpenStudy (alekos):

have you tried to work it out?

OpenStudy (anonymous):

yeah & I came out with 3. & i have a feeling it's not right

Nnesha (nnesha):

can you do it again careful with signs !!!

OpenStudy (anonymous):

\[x+\frac{5}{6}=\frac{1}{3}\\ x=\frac{1}{3}-\frac{5}{6}\]

OpenStudy (anonymous):

so yeah 3 is not right

OpenStudy (anonymous):

\[\frac{1}{3}-\frac{5}{6}=\frac{2}{6}-\frac{5}{6}=-\frac{3}{6}=-\frac{1}{2}\]

OpenStudy (anonymous):

i see that it is not one of your choices so there must be a typo somewhere, unless it is actually \[\frac{x+5}{6}=\frac{1}{3}\]

OpenStudy (anonymous):

if it i the second one, then multiply by 6 and start with \[x+5=2\]

OpenStudy (anonymous):

in which case \(x=2-5=-3\)

Nnesha (nnesha):

 now you can get right answer